OpenStudy (anonymous):

Write the solution in lowest terms. 1/3 - (-3/4) = ANYONE PLEASE :D

OpenStudy (anonymous):

thats not one of the anwsers on the choices

OpenStudy (anonymous):

4/12 - (-9/12) = 4/12 + 9/12 = 13/12 or 1 1/12

Directrix (directrix):

List the choices, please.